# PC-SRIF: Preconditioned Cholesky-based Square Root Information Filter for Vision-aided Inertial Navigation

### v2

@@ II Information Filters Review and PC-SRIF
Notably, when $m$ dominates $n$, the major computational cost of IF is the construction of $\mathcal{I}^{\oplus}$, with the matrix multiplication requiring $O(2mn^{2})$ FLOPs. In SRIF, the most computationally demanding operation is the QR factorization, also requiring $O(2mn^{2})$ FLOPs. However, because $\mathcal{I}^{\oplus}$ is symmetric, only approximately half of its elements need to be computed. Consequently, IF actually requires only $O(1mn^{2})$ FLOPs, approximately twice as efficient as SRIF.
@@
By employing this approach, PC-SRIF achieves a significant reduction in the complexity [$O(mn^{2}_{\scriptscriptstyle 2})$ FLOPs].
@@ III-A Estimated States
$\mathbf{x}_{state}$ $=\begin{bmatrix}{{\mathbf{b}^{\top}_{g_{k}}}\ {{\mathbf{b}^{\top}_{a_{k}}}}\ {}^{\scriptscriptstyle G}{\mathbf{v}}_{\scriptscriptstyle I_{k}}^{\top}\ {{\mathbf{x}}^{\top}_{slam}}\ t_{sync}\ {{\mathbf{x}}^{\top}_{imu}}\ {{\mathbf{x}}^{\top}_{cam}}}\end{bmatrix}^{\top}$ (15)
@@
where $\{G\}$, $\{I\}$, and $\{C\}$ are the global, IMU, and camera frame respectively, and may be sub-scripted by time steps. $\mathbf{b}_{g_{k}},\mathbf{b}_{a_{k}}$ are the biases of gyroscope and accelerometer at time step $k$. ${}^{\scriptscriptstyle G}{\mathbf{v}}_{\scriptscriptstyle I_{k}}$ is the velocity of the IMU w.r.t $\{G\}$. $\mathbf{x}_{poses}$ is the sliding window of size $l$ consisting of IMU poses w.r.t $\{G\}$, where $\mathbf{q}$ is the orientation quaternion and $\mathbf{p}$ is the position. $\mathbf{x}_{slam}$ contains $s$ visual features represented in the inverse depth parameterization [28], where each of them is w.r.t a camera frame in the sliding window. When a past pose slides out the window, the features w.r.t to its frame are reparameterized to frames of the latest pose. $t_{sync}$ is the time offset between the IMU and camera clock. $\mathbf{x}_{cam}$ consists of camera-IMU extrinsics and camera intrinsics $\zeta\in\mathbf{R}^{4}$, focal lengths and camera center.
@@ IV-A Experiment Setup
The goal of our experiments is to provide a comparative analysis of various filtering-based VINS estimators and validate the theoretical advantages of PC-SRIF on practical platforms with limited computational resources. To this end, we implemented a VINS with three estimator options: KF, SRIF, and PC-SRIF. Except for the estimator, all other modules such as image processing and measurement function evaluation are shared, ensuring a fair comparison among the estimators themselves. In this VINS, the number of poses in the sliding window is set to $l=11$, and at most 15 SLAM features and 35 MSCKF features (as defined in [22]) are processed at each time step. Note that since all these estimators are mathematically equivalent, while they may exhibit different efficiency, the only source of difference in their estimates is numerical error.

### v3

@@ II Information Filters Review and PC-SRIF
Notably, when $m$ dominates $n$, the major computational cost of IF is the construction of $\mathcal{I}^{\oplus}$, with the matrix multiplication requiring $O(2mn^{2})$ FLOPs (the Cholesky factorization takes $O(\frac{1}{3}n^{3})$ FLOPs and thus is not major). In SRIF, the most computationally demanding operation is the QR factorization, also requiring $O(2mn^{2})$ FLOPs. However, because $\mathcal{I}^{\oplus}$ is symmetric, only approximately half of its elements need to be computed. Consequently, IF actually requires only $O(1mn^{2})$ FLOPs, approximately twice as efficient as SRIF.
@@
By employing this approach, PC-SRIF achieves a significant reduction in the complexity [$O(mn^{2}_{\scriptscriptstyle 2})$ FLOPs], while being as stable as SRIF. Note that in above steps, (12) should be directly constructed, instead of forming the worse-conditioned (11) and then applying preconditioners to it. This is because the formation of the information matrix in (11) can result in significant irreversible information loss [28, Chapter 5]. For similar reasons, applying preconditioning to an IF (PC-IF) does not effectively address the conditioning issues, which is another point of using SRIF as our framework.
@@ III-A Estimated States
$\mathbf{x}_{state}$ $=\begin{bmatrix}{{\mathbf{b}^{\top}_{g_{k}}}\ {{\mathbf{b}^{\top}_{a_{k}}}}\ {}^{\scriptscriptstyle G}{\mathbf{v}}_{\scriptscriptstyle I_{k}}^{\top}\ {{\mathbf{x}}^{\top}_{slam}}\ t_{sync}\ {{\mathbf{x}}^{\top}_{poses}}\ {{\mathbf{x}}^{\top}_{cam}}}\end{bmatrix}^{\top}$ (15)
@@
where $\{G\}$, $\{I\}$, and $\{C\}$ are the global, IMU, and camera frame respectively, and may be sub-scripted by time steps. $\mathbf{b}_{g_{k}},\mathbf{b}_{a_{k}}$ are the biases of gyroscope and accelerometer at time step $k$. ${}^{\scriptscriptstyle G}{\mathbf{v}}_{\scriptscriptstyle I_{k}}$ is the velocity of the IMU w.r.t $\{G\}$. $\mathbf{x}_{poses}$ is the sliding window of size $l$ consisting of IMU poses w.r.t $\{G\}$, where $\mathbf{q}$ is the orientation quaternion and $\mathbf{p}$ is the position. $\mathbf{x}_{slam}$ contains $s$ visual features represented in the inverse depth parameterization [29], where each of them is w.r.t a camera frame in the sliding window. When a past pose slides out the window, the features w.r.t to that frame are reparameterized to the latest frame. $t_{sync}$ is the time offset between the IMU and camera clock. $\mathbf{x}_{cam}$ consists of camera-IMU extrinsics and camera intrinsics $\zeta\in\mathbf{R}^{4}$, focal lengths and camera center.
@@ IV-A Experiment Setup
The goal of our experiments is to provide a comparative analysis of various filtering-based VINS estimators and validate the theoretical advantages of PC-SRIF on practical platforms with limited computational resources. To this end, we implemented a VINS with three estimator options: KF, SRIF, and PC-SRIF. Except for the estimator, all other modules such as image processing and measurement function evaluation are shared, ensuring a fair comparison among the estimators themselves. In this VINS, the number of poses in the sliding window is set to $l=11$, and at most 15 SLAM features and 35 MSCKF features (as defined in [22]) are processed at each time step. Under this setup, the maximum measurement dimension $m$ is 995995995, and the state size in the update $n_{\scriptscriptstyle 2}$ is 122. Note that since all these estimators are mathematically equivalent, while they may exhibit different efficiency, the only source of difference in their estimates is numerical error.